# SMC-NCA: Semantic-guided Multi-level Contrast for Semi-supervised Action Segmentation

### v1

@@ Abstract
Semi-supervised action segmentation aims to perform frame-wise classification in long untrimmed videos, where only a fraction of videos in the training set have labels. Recent studies have shown the potential of contrastive learning in unsupervised representation learning using unlabelled data. However, learning the representation of each frame by unsupervised contrastive learning for action segmentation remains an open and challenging problem. In this paper, we propose a novel Semantic-guided Multi-level Contrast scheme with a Neighbourhood-Consistency-Aware unit (SMC-NCA) to extract strong frame-wise representations for semi-supervised action segmentation. Specifically, for representation learning, SMC is firstly used to explore intra- and inter-information variations in a unified and contrastive way, based on dynamic clustering process of the original input, encoded semantic and temporal features. Then, the NCA module, which is responsible for enforcing spatial consistency between neighbourhoods centered at different frames to alleviate over-segmentation issues, works alongside SMC for semi-supervised learning. Our SMC outperforms the other state-of-the-art methods on three benchmarks, offering improvements of up to 17.8 and 12.6 in terms of edit distance and accuracy, respectively. Additionally, the NCA unit results in significant better segmentation performance against the others in the presence of only 5 labelled videos. We also demonstrate the effectiveness of the proposed method on our Parkinson’s Disease Mouse Behaviour (PDMB) dataset. The code and datasets will be made publicly available.
@@
ACTION recognition on trimmed videos has achieve remarkable performance over the past few years [1, 2, 3, 4, 5]. Despite the success of these approaches on trimmed videos with a single action, their ability to handle long videos containing multiple actions with different lengths is limited. Temporal action segmentation aims at temporally detecting and recognising human action segments in long and untrimmed videos [6, 7, 8], which has attracted a lot of attention in recent years. This task relies on fully labelled datasets, leading to significant frame-wise annotation cost. To alleviate this problem, many researchers have started exploring weakly supervised approaches using transcripts [9, 10, 11], action sets [12, 13, 14] or annotated timestamps [15] to reduce the annotation cost whilst maintaining action segmentation performance. However, significant efforts are needed to produce partial labels for supervised learning.
@@
Recently, semi-supervised approaches [16, 17, 18] for this task have attracted increasing attention, with a small percentage of labelled videos in the training set. Iterative-Contrast-Classify (ICC) [16] is the first attempt to explore semi-supervised learning for human action segmentation, which consists of two steps, i.e., unsupervised representation learning based on contrastive learning [19, 20] (Fig. 1(a)) and supervised learning. Constructing both positive and negative sets for contrastive learning mainly depends on the clustering results (i.e., labels generated by clustering) of pre-trained input features. Empirical evidence has shown that the outcome of contrastive learning can be severely affected by clustering errors. Besides, only long-term temporal features are leveraged for similarity contrast, which would also limit the ability of representation learning. Another method for this topic is leveraging the correlation of actions between the labelled and unlabelled videos, and temporal continuity of actions [17]. However, their method only relies on action frequency prior in labelled data to build such correlations, rendering it inadequate for guiding the learning of complex unlabelled videos. Consequently, how to effectively explore the critical relations underlying the labelled and unlabelled videos is still an open and challenging problem.
In this paper, we propose a novel SMC-NCA framework for semi-supervised action segmentation. We first propose a novel Sematic-guided Multi-level Contrast (SMC) strategy to enhance frame-level representation learning by combining semantic and temporal information at the same time. Specifically, we introduce a long-term feature extractor and a semantic feature extractor to yield temporal and semantic information, respectively, as shown in Fig. 1(b). Semantic information emphasises action-specific characteristics in contrast to temporal information that encodes the dependencies between actions. Hence, we argue that leveraging the two types of information can help to facilitate contrastive learning. The two extractors are then jointly optimised by our SMC for fully exploring inter-information (between semantic and temporal information) and intra-information variations. Different from [16], our loss is built on triplet losses [21, 22, 23] where anchor, positive and negative representations need to be created. Therefore, we regard the semantic and temporal information as the anchor and positive representations to form positive pairs (i.e., pairwise boxes connected by solid lines in Fig. 1(b)) while explicitly constructing three kinds of negative pairs (e.g., pairwise boxes connected by dotted lines in the semantic branch) densely sampled from the anchor or negative representations. In particular, the generation of the negative pairs is determined by our proposed dynamic clustering scheme on the original input, encoded semantic and temporal features. In contrast, producing positive pairs relies on the inherent similarity between semantic and temporal information without knowing the clustering results.
@@
Inspired by [16], we then iteratively perform supervised learning for action segmentation and unsupervised contrasting for representation updating. However, there still exist significant over-segmentation errors [24] when we use a small amount of labelled data for training the system. Thus, we further propose a Neighbourhood-Consistency-Aware (NCA) unit to explore the consistency between neighbourhoods (i.e., action segments) centered at different frames. Intuitively, feature distributions within the neighbourhoods of the same label should be spatially close. This NCA module can be combined with the SMC method to deal with the over-segmentation issue, particularly when we conduct learning with a small amount of labelled data.
@@
We propose a novel SMC-NCA framework for semi-supervised temporal action segmentation, where semantic and temporal information are jointly utilised to enhance representation learning.
We propose a novel Semantic-guided Multi-level Contrast (SMC) scheme for unsupervised representation learning, allowing one to fully explore intra- and inter-information variations for learning discriminative frame-wise representations with the support of semantic information. This is achieved by constructing positive pairs and three distinct types of complementary negative pairs to facilitate contrastive learning.
We propose a Neighbourhood-Consistency-Aware (NCA) unit to alleviate the over-segmentation problem in semi-supervised settings by utilising inter-neighbourhood consistency. This enables significant improvements of action segmentation in the case of a low number of labelled videos.
Experimental results demonstrate that our proposed SMC scheme significantly improves the capability of frame-wise representation. Our SMC-NCA framework outperforms state-of-the-art methods across different settings of labelled data (5, 10, 40 and 100) on three challenging datasets and our mouse behaviour dataset.
@@ II-A Temporal Action Segmentation
Temporal action segmentation has been an increasingly popular trend in the domain of video understanding, which involves different levels of supervision. Fully-supervised methods require that each frame in the training videos be labelled. Earlier approaches have attempted to incorporate high-level temporal modelling over frame-wise classifiers. Kuehne et al. [25] utilised Fisher vectors of improved dense trajectories to represent video frames, and modelled each action with a Hidden Markov Model (HMM). Singh et al. [26] used a two-stream network to learn representations of short video chunks, which are then fed into a bi-directional LSTM to capture dependencies between different chunks. However, their approaches are computationally intensive due to sequential prediction. Most recent approaches employ TCN (temporal convolutional network) [6, 27, 28], GCN (graph convolution network) [29, 30] or Transformer [31] to model the long-range temporal dependencies in the videos. In spite of their promising performance, obtaining annotations on fine-grained actions for each frame is costly. Weakly-supervised methods mainly focus on transcript-level (with action order information) [9, 10, 11], set-level (without action order information) [12, 13, 14] and timestamp-level [15] supervisions to reduce the annotation effort. However, necessary supervision information is also needed for each video in the training set. Unsupervised approaches [32, 33, 34, 35] take advantages of clustering algorithms without any supervised signal per video, which still suffers from poor performance compared to other supervised settings.
@@ II-B Unsupervised Contrastive Representation Learning
Recently, contrastive representation learning has been extensively used in different computer vision applications such as image representation learning [36, 19, 37, 38], video representation learning [39, 40, 41, 42, 43, 44, 20], time series representation learning [45, 46, 23] and face recognition [21, 47, 48, 49]. Among these approaches, contrastive loss[19, 36] and triplet loss [21, 23] are two popular loss functions to perform contrasting. The underlying idea behind the former is to pull together representations of augmented samples of the same image or video clips (i.e., positive pair) while pulling apart those of different instances (i.e., negative pair). The latter has the same goal but involves defining a triplet of (anchor, positive and negative pairs), where the positive pairs (i.e., anchor and positive representations) should be close and the negative pairs (i.e., anchor and negative representations) should be far apart. Our proposed framework is based on the triplet loss, which is the first attempt to explore complementary information between temporal and semantic information for semi-supervised temporal action segmentation.
@@ III Proposed Method
The overall framework of our SMC-NCA is illustrated in Fig. 2. For unsupervised representation learning, given the input video features, a long-term feature extractor T and semantic feature extractor S are jointly adopted to extract temporal and semantic information. Our SMC is able to not only learn intra-information variation but also inter-information variation, where the temporal information describes the temporal dependencies between actions while semantic information highlights the action-specific characteristics. For semi-supervised learning, the models T, S and a linear classifier C are firstly integrated to perform supervised learning on a small number of labelled videos, followed by a new round of contrastive learning based on the pseudo-labels provided by C. An additional NCA unit (G) is designed in conjunction with contrastive learning to measure the spatial consistency between different neighbourhoods for the alleviation of over-segmentation errors. Following [16], we perform classification and contrastive learning in an iterative manner.
@@ III-C Semantic-guided Multi-level Contrast
As mentioned earlier, we propose to leverage both temporal and semantic information to our proposed SMC module, where the former describes long-range temporal dependencies in long videos while the latter encodes high-level features at each action/frame. In the SMC module, inter- and intra-information variations are jointly learned to enhance the discriminative power of the temporal model.
Temporal and semantic information generation. To encode the correlation between actions, we adopt the C2F-TCN [16] model to produce a multi-resolution representation (denoted as , where is the minibatch size during training, represents the feature dimension of the generated representation) with temporal information used for unsupervised contrastive learning. In the meanwhile, we employ the multilayer perceptron (MLP) as a semantic feature extractor on the same input video features V to generate a new representation (denoted as ) with high-level semantic information. A MLP offers simplicity and efficiency (we also tested other extractors, such as deeper MLPs and 1D CNNs, see Tab. IV). The model S can be formulated as:
where denotes the input hidden representation from the previous layer (with ) and and are trainable parameters. is an activation function.
Inter-information variation learning. In our contrastive learning framework, we fully explore the complementary information between the temporal and semantic information, based on the encoded representation X and H. To achieve this, we firstly learn the similarity and dissimilarity between the two different representations. Specifically, we firstly sample a fixed number of frames from each training video to form a new temporal representation and semantic representation using the same sampling strategy reported in [16], where is the number of the sampled frames from each video. It is too computationally expensive to measure each frame of every video in a mini-batch. The sampled frames for each video are then combined together by concatenation operation on the dimension to form and .
@@
where represents element-wise product. represents the identity matrix used to select pairwise frames with the same index, and is a scale factor to adjust the correlation between vectors. is the element of matrix .
The advantages of optimising this function are two-folds. Firstly, we consider the temporal and semantic information simultaneously, which is beneficial to improving the ability of representation learning. Secondly, learning of positive pairs (anchor and positive representations) is not dependent on the coarse clustering results, compared to ICC that generates related contrastive instances based on clustering results. In contrast, performs contrastive learning with the guide of the potential similarity between frame-wise temporal and semantic representations. We then compute the overall loss between positive pairs below:
@@
The clustering results on the standard input features cannot be updated, which may lead to accumulation errors caused by incorrect negative pairs. To reduce these errors, we perform the clustering on the sampled input , temporal and semantic features simultaneously for dynamic selection of potential negative pairs. Hence, we obtain and for and , respectively. By combining the three matrices, we generate a matrix that is used to dynamically guide the dissimilarity learning on the negative pairs, defined as:
where each element indicates whether or not the -th and -th frames share the same clustering label.
@@
Afterwards, we can construct a set of dense negative pairs for each frame based on M, and negative pairs must be separate. Hence, similar to Eq. (4), the function to be minimised with the corresponding loss is expressed as:
where denotes the number of all the negative pairs.
@@
Intra-information variation learning. and defined in Eqs. (5) and (9) encourage the temporal model to focus on the inter-information variations. Although they implicitly learn the dissimilarity between frames with the semantic or temporal representation, the discriminability of the learned representations is still weak (this can be further verified in Section IV-C1). Therefore, we further explore the intra-information variations, including the dissimilarities within the semantic and temporal representations. For each frame of semantic representations (anchor), we choose all the other frames with different clustering labels to form a set of dense intra-information negative pairs. The overall loss for measuring the dissimilarity between negative pairs with anchor representation can be computed as:
Additionally, we consider the dissimilarity between frames of temporal representations (positive), and the corresponding loss is defined as follows:
@@
Finally, integrating inter-information losses (i.e., Eqs. (5) and (9)) with the intra-information losses in Eqs. (11) and (13), our proposed sematic-guided multi-level contrast loss is formulated as:

### v3

# SMC-NCA: Semantic-guided Multi-level Contrast for Semi-supervised Temporal Action Segmentation

@@ Abstract
Semi-supervised temporal action segmentation (SS-TAS) aims to perform frame-wise classification in long untrimmed videos, where only a fraction of videos in the training set have labels. Recent studies have shown the potential of contrastive learning in unsupervised representation learning using unlabelled data. However, learning the representation of each frame by unsupervised contrastive learning for action segmentation remains an open and challenging problem. In this paper, we propose a novel Semantic-guided Multi-level Contrast scheme with a Neighbourhood-Consistency-Aware unit (SMC-NCA) to extract strong frame-wise representations for SS-TAS. Specifically, for representation learning, SMC is first used to explore intra- and inter-information variations in a unified and contrastive way, based on action-specific semantic information and temporal information highlighting relations between actions. Then, the NCA module, which is responsible for enforcing spatial consistency between neighbourhoods centered at different frames to alleviate over-segmentation issues, works alongside SMC for semi-supervised learning (SSL). Our SMC outperforms the other state-of-the-art methods on three benchmarks, offering improvements of up to 17.8 and 12.6 in terms of Edit distance and accuracy, respectively. Additionally, the NCA unit results in significantly better segmentation performance in the presence of only 5 labelled videos. We also demonstrate the generalizability and effectiveness of the proposed method on our Parkinson’s Disease Mouse Behaviour (PDMB) dataset. The code is publicly available at https://github.com/FeixiangZhou/SMC-NCA.
@@
ACTION recognition on trimmed videos has achieved remarkable performance over the past few years [1, 2, 3, 4, 5]. Despite the success of these approaches on trimmed videos with a single action, their ability to handle long videos containing multiple actions with different lengths is limited. Temporal action segmentation (TAS) aims at temporally detecting and recognising human action segments in a long untrimmed video [6, 7, 8], which has attracted a lot of attention in recent years. Different from temporal action detection (TAD) [9] that detects the temporal boundaries (i.e., start and end) of action instances and predicts corresponding action categories simultaneously using temporally sparse action annotations, the goal of TAS is to produce frame-wise dense action labels. Besides, TAD operates with general videos from everyday life, such as THUMOS14 [10], while TAS targets procedural activity datasets. It is a key task for analysing and understanding human activities in complex long videos, and has been widely applied in fields such as video surveillance [11], video summarization [12] and autonomous vehicles [13]. The temporal relations among sequential human actions are crucial in TAS as they determine the sequence and timing of each action. Recent research [14] reveals that semantic features based on label-text prompts can enhance action segmentation. Despite the impressive performance of the previous approaches [15, 16, 17, 14], they rely on fully labelled videos where the start and end frames of each action are annotated, leading to substantial frame-wise annotation costs. To alleviate this problem, many researchers have started exploring weakly supervised approaches using transcripts [18, 19, 20], action sets [21, 22, 23] or annotated timestamps [24] to reduce the annotation cost whilst maintaining action segmentation performance. However, significant efforts are needed to produce partial labels for supervised learning. Recently, semi-supervised approaches [25, 26, 27] for this task have attracted increasing attention, with a small percentage of labelled videos in the training set.
@@
SS-TAS is a non-trivial task as it often involves processing long untrimmed videos, presenting unique challenges compared to trimmed sequences. It can be difficult to effectively utilising the complex information contained within these unlabelled videos while ensuring accurate segmentation. ICC [25] is the first attempt to explore semi-supervised learning for TAS by introducing a contrast-classify framework, which consists of two steps, i.e., unsupervised contrastive representation learning for frame-wise representation enhancement, as well as semi-supervised learning for action classification. As shown in Fig. 1(a), the contrastive learning approach in ICC primarily focuses on variations within temporal information containing relations between actions, which may ignore the potential role of action-specific semantic characteristics. Meanwhile, constructing both positive and negative sets for contrastive learning mainly depends on the clustering results (i.e., labels generated by clustering) of pre-trained input features. Empirical evidence has shown that the outcome of contrastive learning can be severely affected by clustering errors (see Fig. 2 and Tab. I (the first set of experiments)). Another method for this topic is leveraging the correlation of actions between the labelled and unlabelled videos, and temporal continuity of actions [26]. However, their method only relies on action frequency prior in labelled data to build such correlations, rendering it inadequate for guiding the learning of complex unlabelled videos. Consequently, how to effectively leverage unlabelled, long untrimmed videos is still an open and challenging problem.
In this paper, we propose SMC-NCA, a novel temporal-semantic contrast framework for SS-TAS. Our work is also built on the contrast-classify architecture, but unlike ICC, which contrasts relevant frames within temporal information alone, our SMC incorporates both temporal information (i.e., temporal relations between actions) and semantic information (i.e., action-specific characteristics) into a unified contrastive learning framework, effectively reducing clustering errors and promoting frame-wise representation learning. We further design a NCA unit to handle continuously fluctuating action category predictions, a.k.a. over-segmentation issues, leading to better segmentation performance.
Inspired by the alignment of different augmented views of the same image/video [28, 29, 30], as well as that of multiple modalities [31, 14] (e.g., vision and language), we incorporate a simple semantic feature extractor to extract semantic information, which complements the temporal information for cross-information contrastive representation learning. Semantic information emphasises action-specific characteristics [14] in contrast to temporal information that encodes the dependencies between actions. Hence, we argue that leveraging the two types of information can help to facilitate representation learning (see Fig.1(c)). As shown in Fig. 1(b), temporal and semantic branches are then jointly optimised by our SMC for fully exploring inter- and intra-information variations. Concretely, unlike [25] that recognises positive pairs based on the clustering results, SMC creates them (i.e., pairwise boxes connected by solid lines in Fig. 1(b)) by directly aligning temporal-semantic information, relying on the inherent temporal-semantic consistency of each frame. Meanwhile, we explicitly construct three types of complementary negative pairs to enlarge inter-class dissimilarity, where the generation process is determined by our proposed dynamic clustering algorithm. With the proposed temporal-semantic contrast, our SMC pays attention to action-specific semantic features and temporal relations simultaneously, effectively reducing the impact of clustering errors, thus achieving enhanced representations (see Tab. I).
@@
Benefiting from strong representations learned by SMC, our approach offers a significant improvement on the semi-supervised frame-wise accuracy. However, there still exist significant over-segmentation errors [32] when we use a small amount of labelled data for training the system, resulting in low-quality segmentation results, i.e., segmental F1 score and Edit distance (see Tab. V). This is because the existing contrast-classify framework excessively emphasises frame-wise classification performance by carefully performing inter-frame contrast to enhance the discriminability of each frame’s representation, thus potentially overlooking the consistency of categories within a particular action segment. To this end, we propose NCA to encourage the consistency between neighbourhoods (i.e., action segments) centered at different frames. Intuitively, feature distributions within the neighbourhoods of frames with the same label should be spatially close. NCA can be combined with SMC to deal with over-segmentation issues, particularly when we conduct supervised learning with a small amount of labelled data.
@@
We propose SMC-NCA, a novel temporal-semantic contrast framework tailored for SS-TAS, where temporal information with relations between actions, and semantic information with action-specific attributes are jointly utilised to enhance representation learning.
We propose SMC for unsupervised representation learning, allowing one to fully explore intra- and inter-information variations for learning discriminative frame-wise representations with the support of semantic information.
We propose NCA to alleviate over-segmentation problems in semi-supervised settings by enforcing inter-neighbourhood consistency. This enables significant improvement in the segmentation quality while maintaining good frame-wise accuracy.
Experimental results demonstrate that our SMC-NCA outperforms state-of-the-art methods across different settings of labelled data (5, 10, 40 and 100) on three challenging datasets.
We also introduce a new Parkinson’s Disease Mouse Behaviour (PDMB) dataset, which provides a valuable resource for the research community to investigate the behavioural correlations of mice with Parkinson’s Disease. Experimental results show that our method is generalizable and effective on different kinds of datasets beyond human actions.
@@ II-A Temporal Action Segmentation
Temporal action segmentation (TAS) has been an increasingly popular trend in the domain of video understanding, which involves different levels of supervision. Fully-supervised methods require that each frame in the training videos be labelled. Earlier approaches have attempted to incorporate high-level temporal modelling over frame-wise classifiers. Kuehne et al. [34] utilised Fisher vectors of improved dense trajectories to represent video frames, and modelled each action with a Hidden Markov Model (HMM). Singh et al. [35] used a two-stream network to learn representations of short video chunks, which are then fed into a bi-directional LSTM to capture dependencies between different chunks. However, their approaches are computationally intensive due to sequential prediction. Most recent approaches employ TCN (temporal convolutional network) [6, 15, 33], GCN (graph convolution network) [17, 36] or Transformer [16] to model the long-range temporal dependencies in the videos. In spite of their promising performance, obtaining annotations on fine-grained actions for each frame is costly. Weakly-supervised methods mainly focus on transcript-level (with action order information) [18, 19, 20], set-level (without action order information) [21, 22, 23] and timestamp-level [24] supervisions to reduce the annotation effort. However, necessary supervision information is also needed for each video in the training set. Unsupervised approaches [37, 38, 39, 40] take advantage of clustering algorithms without any supervised signal per video, which still suffer from poor performance compared to other supervised settings.
@@ II-B Unsupervised Contrastive Representation Learning
Visual representation plays a crucial role in various computer vision tasks, including temporal action segmentation [35, 16], person retrieval [41, 42] and event detection [43]. It can be learned in different ways, including fully supervised, semi-supervised and unsupervised methods. For instance, Shi et al. [42] proposed to mine various human attributes and model relations between them to learn precise attribute representations. Shen et al. [43] explored multi-modal information fusion to enrich video content representations.
Recently, contrastive representation learning has been extensively used in different computer vision applications such as image representation learning [44, 28, 45, 46], video representation learning [29, 47, 48, 49, 50, 30], time series representation learning [51, 52, 53], face recognition [54, 55, 56]. Among these approaches, contrastive loss[28, 44] and triplet loss [54, 53] are two popular loss functions to perform contrasting. The underlying idea behind the former is to pull together representations of augmented samples of the same image or video clips (i.e., positive pair) while pulling apart those of different instances (i.e., negative pair). The latter has the same goal but involves defining a triplet of (anchor, positive and negative pairs), where the positive pairs (i.e., anchor and positive representations) should be close and the negative pairs (i.e., anchor and negative representations) should be far apart.
The idea of visual-semantic alignment for representation enhancement has been explored in various fields such as few-shot learning [57, 58], meta learning [59] and out-of-distribution detection [60]. Different from most methods that operate within a supervised setting where label information is utilised to guide feature/image alignment, our SMC performs temporal-semantic contrast in a unsupervised way to fully utilise unlabelled videos. Furthermore, it is based on the triplet loss [53], which effectively reduces the dependence on noisy clustering labels, and leverages complementary information between temporal and semantic information compared with other methods.
II-C Semi-supervised Learning
Semi-supervised learning (SSL) is a good solution to reduce the cost of data labelling, which leverages a few labelled samples and a large number of unlabeled samples to train the model. The current semi-supervised methods are mainly divided into prediction-based and representation-based approaches [61]. Prediction-based approaches typically involve regularizing the classifier’s prediction of unlabelled data, where two common paradigms for this purpose are pseudo-labelling [62, 63, 64] and consistency regularization [65, 66, 67]. Pseudo-labelling based methods generate pseudo labels for unlabeled data through pre-trained models and use these pseudo labels to further optimise the model. Consistency based methods encourage the model to make similar predictions on images obtained by applying different data augmentation techniques to the same image. For example, FixMatch [65] uses the model to generate pseudo labels for weakly augmented unlabeled images and only samples with highly-confident pseudo-labels will be used for training. To adapt FixMatch to video-based action recognition, Wu et al. [68] designed a neighbor-guided consistent and contrastive learning framework to reuse lowly-confident samples and learn more discriminative feature representations. However, prediction-based approaches are less effective compared to representation-based approaches in large-scale scenarios [61].
Representation-based methods focus more on the representations encoded by deep learning networks. Recent studies exploit the success of self-supervised learning [44, 28] in learning representations from unlabeled data to train large-scale semi-supervised models [69, 70]. PAWS [71] proposed a single-stage training scheme that combines supervised and self-supervised learning for SSL. Although these methods have shown promise, it remains challenging to directly apply data augmentation strategies, such as flipping, rotation, and transformation, used in the image or video domain to TAS as the inputs are pre-computed feature vectors [26]. Inspired by [25], we propose a novel temporal-semantic contrast framework for representation enhancement, which shows strong frame-wise representations and segmentation performance.
@@ III Proposed Method
The overall framework of our SMC-NCA is illustrated in Fig. 3. For unsupervised representation learning, given the input video features, a temporal feature extractor T and semantic feature extractor S are jointly adopted to extract temporal and semantic information. Our SMC learns not only intra-information variation but also inter-information variation, where the temporal information describes the temporal dependencies between actions while semantic information highlights the action-specific characteristics. For semi-supervised learning, the models T, S and a linear classifier C are firstly integrated to perform supervised learning on a small number of labelled videos, followed by a new round of contrastive learning based on the pseudo-labels provided by C. An additional NCA unit (G) is designed in conjunction with contrastive learning to measure the spatial consistency between different neighbourhoods for the alleviation of over-segmentation errors. Following [25], we iteratively perform classification and contrastive learning. We have not altered the structure of the backbone C2F-TCN [25]; instead, we introduce two additional modules, S and G, to facilitate unsupervised and semi-supervised learning. These two modules seamlessly integrate with the backbone and are exclusively utilised during model training to improve the semi-supervised segmentation performance. They are not utilised during the inference phase.
@@ III-C Semantic-guided Multi-level Contrast
As mentioned earlier, we propose to leverage both temporal and semantic information to our proposed SMC scheme, where inter- and intra-information variations are jointly learned to enhance frame-wise representations.
Temporal and semantic information generation. To encode the correlation between actions, we adopt the C2F-TCN model to produce the multi-resolution representation and combine representations of different resolutions into a new representation using the same temporal up-sampling approach in [25]. Similar to [25], the combined representation with temporal information is used for unsupervised contrastive learning, which is denoted as , where is the mini-batch size during training and represents the feature dimension.
In ICC, only temporal information is utilised to perform inter-frame contrastive learning. In contrast, we propose to incorporate semantic information alongside temporal information for cross-information contrast. This decision is motivated by the significance of semantic features in capturing action-specific details, a factor that has been proven to be effective in promoting temporal action segmentation [14]. Different from extracting semantic features from action- or class-specific text information [14], we implicitly extract such features directly from the raw input features without the need for complex text encoding models. Specifically, we employ the multilayer perceptron (MLP) as a semantic feature extractor on the input video features V to generate a new representation (denoted as ) with high-level semantic information. A MLP offers simplicity and efficiency (we also test other extractors, such as deeper MLPs and 1D CNNs, see Tab. III). The model S can be formulated as:
where denotes the input hidden representation from the previous layer (with ) and and are trainable parameters. is an activation function. Temporal and semantic information complement each other in promoting representation learning as they capture different aspects of actions. Next, we will elaborate on how to incorporate these two types of information into a unified contrastive learning framework.
Inter-information variation learning. Under the guidance of clustering labels, ICC reduces the intra-class distance by bringing frames belonging to the same label closer together. However, it is susceptible to clustering errors. In practice, the distribution of data may be complex and uneven, leading the clustering algorithm to assign similar frames to different clusters or dissimilar frames to the same cluster. This clustering error can cause ICC to learn incorrect similarity information during training, thus affecting its performance. Additionally, clustering errors may result in overfitting to samples with noisy labels, thereby reducing its generalization ability.
With the support of semantic information, we directly construct inter-information positive pairs without relying on noisy clustering labels. To be more specific, we firstly sample a fixed number of frames from each training video to form a new temporal representation and semantic representation using the same sampling strategy reported in [25], where is the number of the sampled frames from each video. This is because it is too computationally expensive to measure each frame of every video in a mini-batch. The sampled frames for each video are then combined by concatenation operation on the dimension to form and .
@@
where represents the element-wise product. represents the identity matrix used to select pairwise frames with the same index, and is a scale factor to adjust the correlation between vectors. is the element of matrix .
The advantages of optimising this function are twofold. Firstly, we consider the temporal and semantic information simultaneously, which is beneficial to improving the ability of representation learning. Secondly, learning of positive pairs (anchor and positive representations) is not dependent on the coarse clustering results, compared to ICC which generates related contrastive instances based on clustering results. In contrast, performs contrastive learning with the guide of the potential consistency between frame-wise temporal and semantic representations. We then compute the inter-information similarity loss between positive pairs below:
where promotes intra-class similarity by considering the one-to-one correspondence between temporal and semantic information, where each frame located at the same position belongs to the same category.
@@
where indicates the similarity between all sampled frames in a mini-batch. When the clustering labels between the -th and -th frames are identical, the corresponding element in the matrix is set to 1; otherwise, it is set to 0.
@@
The clustering results on the standard input features cannot be updated during training [25], which may lead to accumulation errors caused by incorrect negative pairs. To reduce these errors, we perform the clustering on the sampled input , temporal and semantic features simultaneously for dynamic selection of potential negative pairs. Hence, we obtain and for and , respectively. By combining the three matrices, we generate a matrix that is used to dynamically guide the dissimilarity learning on the negative pairs, defined as:
where each element represents whether the -th and -th frames share the same clustering category. 1 indicates frames from different categories while 0 indicates frames from the same category. The combination of these three matrices can help improve the quality of negative pairs.
@@
Afterwards, we can construct a set of dense negative pairs for each frame based on M, and negative pairs must be separate. Hence, similar to Eq. (4), the function to be minimised with the corresponding inter-information dissimilarity loss is expressed as:
where is utilised to select pairwise frames between temporal and semantic information to construct inter-information negative pairs. denotes the number of all the negative pairs. facilitates the separation of inter-information different categories.
@@
Intra-information variation learning. and defined in Eqs. (5) and (9) encourage the temporal model to focus on inter-information variations (i.e., similarity and dissimilarity). Although they implicitly learn the dissimilarity between frames with the semantic or temporal representation, the discriminability of the learned representations is still weak as the inherent discriminative cues within semantic or temporal information are not exploited (this can be verified in Section IV-C1).
Therefore, we further explore the intra-information variations, including the dissimilarities within the semantic or temporal representation. In detail, for each frame of semantic representation (anchor), we choose all the other frames with different clustering labels to form a set of dense intra-information negative pairs. The intra-information dissimilarity loss based on semantic representation can be computed as:
where is the sigmoid function. , and are defined in Eqs. (7), (4) and (9), respectively. Here is used to select pairwise frames with distinct categories from semantic representation . encourages the separation of different categories within semantic information. Additionally, we consider the dissimilarity between frames of temporal representation (positive), and the corresponding loss is defined as follows:
@@
By integrating inter-information losses (i.e., Eqs. (5) and (9)) with the intra-information losses in Eqs. (11) and (13), our proposed sematic-guided multi-level contrast loss is formulated as:
Notably, by introducing semantic information to the temporal-semantic contrast framework, our method mines more useful information in a large number of unlabelled untrimmed videos. Semantic information plays a multifaceted role in our method: (1) It contributes to the construction of inter-information positive pairs by aligning temporal and semantic information directly, thereby enlarging the intra-class similarity between the two types of information. (2) It also participates in constructing inter-information negative pairs, aiming to enlarge inter-class dissimilarity between the two types of information. (3) It is involved in constructing intra-information negative pairs, further enlarging the intra-class dissimilarity within semantic information. In summary, with the support of semantic information, our SMC can learn representations that capture both the temporal dynamics and semantic nuances of actions, thereby improving the discriminability and robustness of these representations.